A) In a police speed trap a straight length 𝐴𝐵, where 𝐴𝐵 = 186 m, is kept under surveillance. An ambulance from Chreso University passes point 𝐴 at 26.5 ms −1 and accelerating at a constant rate of 1.5 ms−2 .

i) What is the speed of the ambulance when it passes point 𝐵?

ii) What is the time taken to cover this distance?


B) If the speed limit is 70 mph, determine whether the speed limit has been broken?

i) Based on the average speed over the whole distance or

ii) Based on its speed at the moment it passes C, mid-point of 𝐴𝐵. (Take 1 km = 0.622 miles)


1
Expert's answer
2021-09-29T17:50:45-0400

a) i)


"V^2-26.5^2=2(1.5)(186)\\V=35.5\frac{m}{s}"

ii)


"t=\frac{35.5-26.5}{1.5}=6\ s"

b) i) No


"v=0.5(35.5+26.5)=31\frac{m}{s}=69\ mph"

ii) Yes


"u^2-26.5^2=2(0.5)(1.5)(186)\\u=31.3\frac{m}{s}=70.1\ mph"